Given a right triangle with a leg length of 4cm and a hypotenuse lengh of 8cm, find the length of the other leg?
Can anyone help? I keep working out the problem but I am not coming up with any answers provided. Thanks!

Answer by CharlesG2(834)

"Given a right triangle with a leg length of 4cm and a hypotenuse lengh of 8cm, find the length of the other leg?
Can anyone help? I keep working out the problem but I am not coming up with any answers provided. Thanks!"
Pythagorean Theorem: a^2 + b^2 = c^2, where a and b are legs of a right triangle, and c is the hypotenuse.
c^2 - a^2 = b^2
8^2 - 4^2 = b^2
64 - 16 = b^2
48 = b^2
16 * 3 = 48 = b^2
4sqrt(3) = b